We want to share an important update with our Betfair Community users.
Over the coming months, the Betfair Community platform will be winding down. The final active day for the platform will be December 21st. From December 22nd, users will no longer be able to access the platform.
Your discussions, insights, engagement, and, most importantly, feedback have helped shape the platform and guided innovation. Your feedback continues to be important to us, and we will continue to listen via our dedicated feedback channels, as well as through product testing and innovations.
To share your feedback and experience, visit: https://support.betfair.com/app/home
|
By:
(5/6)^9 = 0.194
|
|
By:
That is the probability .^
|
|
By:
If P=a/b then odds against are (b-a)/a.
EG P= 1/3 odds against = (3-1)/1 |
|
By:
I make it odds of 9/5 in my head
odds against 4 appearing with 1 dice are 5/1, therefore with 9 dice 5/9 therefore odds AGAINST must be 9/5, shirley! |
|
By:
I ask as im am going to the casino on sunday and am doing the 6 line and if I win only once in 9 goes I will be up. so I was thinking what Is the chance of winning or losing in this scenario?
|
|
By:
I should let you know its single zero roulette
|
|
By:
kenny, you`re adding the probabilities, rather than multiplying them. If you have a 5/6 and 5/6 double, you don`t get paid at 10/6.
With six dice thrown on average there should be one `four` with every throw. With more than six dice thrown, you will be getting more than one four, on average, meaning a four is more likely to appear than not to appear, making it a mathematical odds-on shot. So, it can`t be 9/5. Andriy has it right, rounded to 1.2 or 1/5 |
|
By:
thanks timbuctooth I thought the other answers looked a little wrong
|
|
By:
Incidentally there is no difference in these odds for one die thrown 9 times and 9 dice thrown simultaneously.
|
|
By:
1.166666666666 x 9 equates to almost exactly 3/1
|
|
By:
1.16 x 9 = 10.44
What has that got to do with OP's question? |
|
By:
Andriy is correct
0.194 = 4.15/1 |
|
By:
^
Andriy did do a sum correctly, but didn't answer the question correctly. Pumphol didn't even do his bizarre sum correctly. You are the first person to actually answer the question. |
|
By:
The odds are 104/25 and the probability is (5/6)^9.
|
|
By:
dave1357 20 Apr 16 13:13
1.16 x 9 = 10.44 What has that got to do with OP's question? Well I was doing 1.166 x 1.166 ( 9 times ) to get the 3/1 Now that may be the wrong formula & even wrong answer but I would have thought you would have at least seen where I was going with the 1.16 & 9 ( the nine equating to nine dice or nine throws ) so the 1.16 x 9 equating to 10.44 is not what I meant especially when I said I made it around 3/1, so your dave1357 20 Apr 16 14:57 Pumphol didn't even do his bizarre sum correctly. My formula or lack of it was wrong but you wrote 10.44 not me, but thanks anyway for being so condescending. |
|
By:
so is it more or less than 80% chance a 4 would not come up = about 5-1 ?
|
|
By:
No, it's a 19.38% chance that a 4 will not come up
|
|
By:
0.833333⁹ = 0.833333×0.833333×0.833333×0.833333×0.833333×0.833333×0.833333×0.833333×0.833333 = 0.19380600176481308
|
|
By:
|
|
By:
thanks evee that sounds like 80% a 4 would come up { is that 4-1 or 5-1
|
|
By:
4.15/1
|
|
By:
Surely that`s ^^^ NOT to come up?
|
|
By:
Getting ridiculous. Are you taking the p.ss AM?
Yes, 4.15/1 that a 4 will NOT come up. |
|
By:
ha ha ha ha
|
|
By:
thanks eeve. seams like about 4.15ish - 1. correct?
|
|
By:
?
|
|
By:
'Ow did you get on that day back in April at the casino, AM?
|
|
By:
ist like poker, you have 54 numbers 9 number 4 so take the 9 from the 54 =45/9hence 5/1
|
|
By:
ist like poker, you have 54 numbers 9 number 4 so take the 9 from the 54 =45/9hence 5/1
Well that clears it up |